# Berry-Esséen bound for the Moment Estimation of the fractional Ornstein–Uhlenbeck model under fixed step size discrete observations

### v1

@@ 1 Introduction and main results
$\begin{split}d_{Kol}\left(\sqrt{n}(\hat{\theta}_{n}-\theta),\mathcal{N}\right)\leq C_{\theta,H,h}\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}),\end{array}\right.\end{split}$ (3)
@@
It is clear that the four Gaussian noises from Example 1.2 to Example 1.5 satisfy Hypothesis 1.6, see [9, 8]. Now we give our second contribution of the present paper.
@@
If the driving noise satisfies Hypothesis 1.6 with $H\in(0,\frac{1}{2})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
@@
Clearly, both Example 1.2 and Example 1.5 satisfy Hypothesis 1.8. An additional requirement $H^{\prime}\in(\frac{1}{2},1)$ for both Example 1.4 and Example 1.3 makes Hypothesis 1.8 hold.
@@
Assume that both the Ornstein-Uhlenbeck model $\{Z_{t}:t\geq 0\}$ and the estimator $\hat{\theta}_{n}$ are given as in Theorem 1.7. If the driving noise satisfies Hypothesis 1.8 with $H\in(\frac{1}{2},\frac{3}{4})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
@@ 2 Preliminaries
[…]
@@
$d_{Kol}(\sqrt{T}(f(X)-\theta),\,\xi)\leq C\times\left(d_{Kol}(\sqrt{T}(X-\mathbb{E}[X]),\eta)+\sqrt{T}\left|\mathbb{E}[X]-g(\theta)\right|+\frac{1}{\sqrt{T}}\right),$
@@
Substituting this result into (17), we obtain the desired result. ∎
@@ 3.1 The second moment and cumulants of the random variable $W_{n}$
$\begin{split}\left|\mathbb{E}(W_{n}^{2})-\sigma_{B}^{2}\right|\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}),\end{array}\right.\end{split}$ (21)
@@
$\begin{split}S_{1}&\leq\frac{C}{n}\left[\sum_{j,l=1}^{n}e^{-\theta h(l+j)}+\sum_{j,l=1}^{n}e^{-\theta jh}\left(1+l\right)^{2H-2}\right]\\ &\leq\frac{C}{n}\left[\int_{1}^{\infty}\int_{1}^{\infty}e^{-\theta h(x+y)}\mathrm{d}x\mathrm{d}y+\int_{1}^{\infty}e^{-\theta hx}\mathrm{d}x\int_{1}^{n}y^{2H-2}\mathrm{d}y\right]\\ &\leq\frac{C}{n}\left[1+n^{(2H-1)\vee 0}\right]\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{2(1-H)}},&\quad\text{if }H\in(\frac{1}{2},\,1).\end{array}\right.\end{split}$ (33)
@@
And then, making the change of variable $k=j-l$ yields
@@
$\begin{split}\frac{4}{n}\sum_{k=1}^{n}k\rho_{0}^{2}(kh)&\leq\frac{C}{n}\sum_{k=1}^{n}(1+k)^{2(2H-2)+1}\leq\frac{C}{n}\int_{n}^{\infty}x^{2(2H-2)+1}\mathrm{d}x\\ &\leq Cn^{(4H-2)\vee 0-1}=C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{1-2H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}).\end{array}\right.\end{split}$ (38)
@@
$\begin{split}S_{2}\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{1-2H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}).\end{array}\right.\end{split}$ (39)
Substituting the estimations (33) and (39) into (28) with the fact that $32(H-1)<4H-$ 3 if $H\in(\frac{1}{2},\,\frac{3}{4})$, we obtain the desired result. ∎
@@
$\begin{split}\max\left\{\left|k_{3}(W_{n})\right|,k_{4}(W_{n})\right\}\leq C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{2}{3}],\\ n^{\frac{3}{2}(4H-3)},&\quad\text{if }H\in(\frac{2}{3},\,\frac{3}{4}).\end{array}\right.\end{split}$ (40)
Proof.
@@
$\begin{split}k_{3}(\overline{W}_{n})&\leq\frac{C}{\sqrt{n}}\left(\sum_{|k|<n}\left|\rho_{0}(k)\right|^{\frac{3}{2}}\right)^{2}\leq\frac{C}{\sqrt{n}}\left(\sum_{k=0}^{n-1}(1+k)^{\frac{3}{2}(2H-2)}\right)^{2}\\ &\leq Cn^{(6H-4)\vee 0-\frac{1}{2}}=C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{2}{3}],\\ n^{\frac{3}{2}(4H-3)},&\quad\text{if }H\in(\frac{2}{3},\,\frac{3}{4}),\end{array}\right.\end{split}$ (44)
$\begin{split}k_{4}(\overline{W}_{n})&\leq\frac{C}{n}\left(\sum_{|k|<n}\left|\rho_{0}(k)\right|^{\frac{4}{3}}\right)^{3}\leq\frac{C}{n}\left(\sum_{k=0}^{n-1}(1+k)^{\frac{4}{3}(2H-2)}\right)^{3}\\ &\leq Cn^{(8H-5)\vee 0-1}=C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ n^{2(4H-3)},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}).\end{array}\right.\end{split}$ (45)
@@
$\begin{split}\sup_{j}\left|\left|R_{j}\right|\right|_{L^{2}(\Omega)}<\infty,\end{split}$ (47)
@@
$\begin{split}\sup_{j}\left|\left|\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}\right|\right|_{L^{2}(\Omega)}<\frac{C}{\sqrt{n}},\end{split}$ (48)
@@
$\begin{split}&\left|k_{3}({W}_{n})-k_{3}(\overline{W}_{n})\right|=\left|\mathbb{E}\left(W_{n}^{3}-\overline{W}_{n}^{3}\right)\right|\\ &=\left|\mathbb{E}\left[\left(\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}-\sqrt{n}\left(\mathbb{E}(B_{n})-a\right)\right)\left(W_{n}^{2}+W_{n}\overline{W}_{n}+\overline{W}_{n}^{2}\right)\right]\right|\\ &\leq\left[\left|\left|\frac{1}{\sqrt{n}}\sum_{j=1}^{n}e^{-\theta jh}R_{jh}\right|\right|_{L^{2}(\Omega)}+\left|\left|\sqrt{n}\left(\mathbb{E}(B_{n})-a\right)\right|\right|_{L^{2}(\Omega)}\right]\left|\left|W_{n}^{2}+W_{n}\overline{W}_{n}+\overline{W}_{n}^{2}\right|\right|_{L^{2}(\Omega)}\\ &\leq\frac{C}{\sqrt{n}},\end{split}$ (49)
where in the last inequality we also have used the estimation (48) and Propositions 3.1, 3.2. A similar method yields
@@
Combining the estimations (44), (45), (49), (50), we can obtain the desired result. ∎
@@ 3.2 Berry-Esséen type upper bound for the moment estimator of fOU process
$\begin{split}d_{Kol}(W_{n},\varpi_{n})&\leq d_{TV}(W_{n},\varpi_{n})\leq C\max\left\{k_{3}(W_{n}),k_{4}(W_{n})\right\}\\ &\leq C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{2}{3}],\\ n^{\frac{3}{2}(4H-3)},&\quad\text{if }H\in(\frac{2}{3},\,\frac{3}{4}),\end{array}\right.\end{split}$ (54)
@@
$\begin{split}d_{Kol}(\varpi_{n},\varpi)\leq\frac{2}{\sigma_{n}^{2}\vee\sigma_{B}^{2}}\left|\sigma_{n}^{2}-\sigma_{B}^{2}\right|\leq C\times\left\{\begin{array}[]{ll}\frac{1}{n},&\quad\text{if }H\in(0,\,\frac{1}{2}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{1}{2},\,\frac{3}{4}).\end{array}\right.\end{split}$ (55)
@@
$\begin{split}(d_{Kol}(\sqrt{n}(B_{n}-\mathbb{E}[B_{n}]),\varpi)\leq C\times\left\{\begin{array}[]{ll}\frac{1}{\sqrt{n}},&\quad\text{if }H\in(0,\,\frac{5}{8}],\\ \frac{1}{n^{3-4H}},&\quad\text{if }H\in(\frac{5}{8},\,\frac{3}{4}).\end{array}\right.\end{split}$ (56)
@@ 4.1 The second moment and cumulants of the random variable $\widetilde{W}_{n}$
$\widetilde{W}_{n}=\sqrt{n}\left(A_{n}-\mathbb{E}(A_{n})\right)=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\left(Z_{jh}^{2}-\mathbb{E}\left(Z_{jh}^{2}\right)\right).$ (1)
@@
Next, we will estimate the upper bound of $D_{1}$ and $D_{2}$, respectively. The estimation (6) in Lemma 5.2, along with the symmetry of $j,l$ and the change variable $k=l-j$ imply that
@@
Consequently, substituting inequality (5), (6) into equation (4) obtains the desired result (3). In summary, this completes the proof. ∎
@@
This complete the proof. ∎
@@
This completes the proof. ∎
@@ 4.2 Proofs of main Theorems
Next, using arguments analogous to those in the estimation of (56) combined with Propositions 4.1, 4.2, 4.3, we derive that for $H\in(0,\frac{1}{2})$,
@@
Proof of Theorem 1.9. The distinction between Theorem 1.7 and Theorem 1.9 lies in improving Hypothesis 1.6 with $H\in(0,\frac{1}{2})$ to Hypothesis 1.8 with $H\in(\frac{1}{2},\frac{3}{4})$, thereby accommodating broader Gaussian processes such as the bi-fractional Brownian motion and the sub-bifractional Brownian motion. Notice that the estimations in Lemma 5.2 play a key rule in the proof of Theorem 1.7. To establish Theorem 1.9, it suffices to build up a new comparison about covariance functions ${\rho}(t,s)$ and $\tilde{\rho}(t,s)$, which is presented in the following Proposition.
@@
Let ${\rho}(t,s)=\mathbb{E}(X_{t}X_{s})$, $\tilde{\rho}(t,s)=\mathbb{E}(Z_{t}Z_{s})$ be the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.8 with $H\in(\frac{1}{2},1)$. Then there exists a constant $C\geq 0$ independent of $T$ such that for any $0\leq s\leq t\leq T$,
@@
At the same time, Hypothesis 1.8 with restriction $H\in(\frac{1}{2},1)$ implies that
@@
under Hypothesis 1.8. Let’s estimate ${\rm II}$ in two parts based on the inequality (43).
Part 1. Making the change of variable $x=v-s$ implies that
@@
where in the last inequality we use Lemma 6 with $H\in(\frac{1}{2},1)$.
Part 2. Suppose that $H^{\prime}\in(\frac{1}{2},1),\,K\in(0,2)$ and $H:=H^{\prime}K\in(\frac{1}{2},1)$. We make the change of variable $x=v-s$ and use the L’Hôpital’s rule to get that
@@ 5 Appendix
by the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.6. Then there exists a positive constant $C$ independent of $T$ such that for any $0\leq s\leq t\leq T$,

### v2

@@ 1 Introduction and main results
[…]
@@
It is clear that the four Gaussian noises from Example 1.3 to Example 1.6 satisfy Hypothesis 1.7, see [8, 9]. Now we give our second contribution of the present paper.
@@
If the driving noise satisfies Hypothesis 1.7 with $H\in(0,\frac{1}{2})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
@@
Clearly, both Example 1.3 and Example 1.6 satisfy Hypothesis 1.9. An additional requirement $H^{\prime}\in(\frac{1}{2},1)$ for both Example 1.5 and Example 1.4 makes Hypothesis 1.9 hold.
@@
Assume that both the Ornstein-Uhlenbeck model $\{Z_{t}:t\geq 0\}$ and the estimator $\hat{\theta}_{n}$ are given as in Theorem 1.8. If the driving noise satisfies Hypothesis 1.9 with $H\in(\frac{1}{2},\frac{3}{4})$, then there exists a positive constant $C_{\theta,H,h}$ independent of $n$ such that when $n$ is sufficiently large,
@@ 2 Preliminaries
$\langle\mathbbm{1}_{[a,b]},\,\mathbbm{1}_{[c,d]}\rangle_{\mathfrak{H}}=\mathbb{E}\left((G_{b}-G_{a})(G_{d}-G_{c})\right),$
@@
$d_{Kol}(\sqrt{T}(f(X)-\theta),\,\xi)\leq C\times\left(d_{Kol}(\sqrt{T}(X-\mathbb{E}[X]),\eta)+\sqrt{T}\left|\mathbb{E}[X]-g(\theta)\right|+\frac{1}{\sqrt{T}}\right),$ (10)
@@
Substituting this result into (18), we obtain the desired result.
@@ 3.1 The second moment and cumulants of the random variable $W_{n}$
[…]
@@
[…]
@@
And then, making the change of variable k=jlk=j-l yields
@@
[…]
@@
[…]
Substituting the estimations (34) and (40) into (29) with the fact that 2(H1)<4H32(H-1)<4H-3 if $H\in(\frac{1}{2},\,\frac{3}{4})$, we obtain the desired result.
@@
[…]
Proof 3.6.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
where in the last inequality we also have used the estimation (49) and Propositions 3.1, 3.3. A similar method yields
@@
Combining the estimations (45), (46), (50), (51), we can obtain the desired result.
@@ 3.2 Berry-Esséen type upper bound for the moment estimator of fOU process
[…]
@@
[…]
@@
[…]
@@ 4.1 The second moment and cumulants of the random variable $\widetilde{W}_{n}$
$\widetilde{W}_{n}:=\sqrt{n}\left(A_{n}-\mathbb{E}(A_{n})\right)=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\left(Z_{jh}^{2}-\mathbb{E}\left(Z_{jh}^{2}\right)\right).$ (1)
@@
Next, we will estimate the upper bound of $D_{1}$ and $D_{2}$, respectively. The estimation (6) in Lemma 5.2, along with the symmetry of $j,l$ and the change variable k=ljk=l-j imply that
@@
Consequently, substituting inequality (5), (6) into equation (4) obtains the desired result (3). In summary, this completes the proof.
@@
This complete the proof.
@@
This completes the proof.
@@ 4.2 Proofs of main Theorems
Next, using arguments analogous to those in the estimation of (57) combined with Propositions 4.1, 4.3, 4.5, we derive that for $H\in(0,\frac{1}{2})$,
@@
Proof of Theorem 1.10. The distinction between Theorem 1.8 and Theorem 1.10 lies in improving Hypothesis 1.7 with $H\in(0,\frac{1}{2})$ to Hypothesis 1.9 with $H\in(\frac{1}{2},\frac{3}{4})$, thereby accommodating broader Gaussian processes such as the bi-fractional Brownian motion and the sub-bifractional Brownian motion. Notice that the estimations in Lemma 5.2 play a key rule in the proof of Theorem 1.8. To establish Theorem 1.10, it suffices to build up a new comparison about covariance functions ${\rho}(t,s)$ and $\tilde{\rho}(t,s)$, which is presented in the following Proposition.
@@
Let ${\rho}(t,s)=\mathbb{E}(X_{t}X_{s})$, $\tilde{\rho}(t,s)=\mathbb{E}(Z_{t}Z_{s})$ be the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.9 with $H\in(\frac{1}{2},1)$. Then there exists a constant $C\geq 0$ independent of $T$ such that for any $0\leq s\leq t\leq T$,
@@
At the same time, Hypothesis 1.9 with restriction $H\in(\frac{1}{2},1)$ implies that
@@
under Hypothesis 1.9. Let’s estimate $\rm II$} in two parts based on the inequality (43).
Part 1. Making the change of variable x=vsx=v-s implies that
@@
where in the last inequality we use Lemma 5.1 with $H\in(\frac{1}{2},1)$.
Part 2. Suppose that $H^{\prime}\in(\frac{1}{2},1),\,K\in(0,2)$ and $H:=H^{\prime}K\in(\frac{1}{2},1)$. We make the change of variable x=vsx=v-s and use the L’Hôpital’s rule to get that
@@ 5 Appendix
by the covariance function of the Ornstein-Uhlenbeck processes $X_{t}$ and $Z_{t}$ driven by fBm $B_{t}^{H}$ and $G_{t}$ satisfying Hypothese 1.7. Then there exists a positive constant $C$ independent of $T$ such that for any $0\leq s\leq t\leq T$,
@@
Next, we provide a detailed proof of the key Lemma 2.1 of this paper.
Proof of Lemma 2.1. Without loss of the generality, we assume $\sigma_{1}=1$ and generalize the function $f(x)=x^{-\frac{1}{\alpha}},\,0<\alpha<2$. Then we have $\xi\sim N(0,1)$ and $\Phi(z)$ is its cumulative distribution function. Thus, we have
$\begin{split}d_{Kol}(\sqrt{T}(f(X)-\theta),\,\xi)=\sup_{z\in\mathbb{R}}\left|P\left({\sqrt{T}}(f(X)-\theta)\leq z\right)-\Phi(z)\right|\end{split}$ (10)
Denote
$\begin{split}A(z):=P\left({\sqrt{T}}(f(X)-\theta)\leq z\right)-\Phi(z).\end{split}$ (11)
Since $X\geq 0$ almost surely and $f(x)=x^{-\frac{1}{\alpha}},\,0<\alpha<2$, we have $f(X)>0$ a.s. Hence, we shall assume $z>-\sqrt{T}\theta$. Otherwise, the standard estimate for a normal random variable $\Phi(-t)\leq\frac{1}{2t},\,\forall t>0$ yields
$|A|=\Phi(z)\leq\Phi\big(-\sqrt{T}\theta\big)\leq\frac{C}{\sqrt{T}}\,.$
When $z>-\sqrt{T}\theta$, we have $\frac{z}{\sqrt{T}}+\theta>0$. Hence, the monotonicity of the function $f(x)=x^{-\frac{1}{\alpha}}$ implies that
$\left\{X^{-\frac{1}{\alpha}}-\theta\leq\frac{z}{\sqrt{T}}\right\}=\left\{X-\theta^{-\alpha}\geq\left(\frac{z}{\sqrt{T}}+\theta\right)^{-\alpha}-\theta^{-\alpha}\right\}.$
Recall that $g(\theta)=\theta^{-\alpha}$. We have when $z>-\sqrt{T}\theta$,
$\begin{split}A(z)&=P\left(X^{-\frac{1}{\alpha}}-\theta\leq\frac{z}{\sqrt{T}}\right)-\Phi(z)\\ &=P\left(X-\theta^{-\alpha}\geq\left(\frac{z}{\sqrt{T}}+\theta\right)^{-\alpha}-\theta^{-\alpha}\right)-\Phi(z)\\ &=P\left(\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-g(\theta)\right)\geq\sqrt{\frac{T}{\sigma_{2}^{2}}}\theta^{-\alpha}\left[\left(\frac{z}{\sqrt{T}\theta}+1\right)^{-\alpha}-1\right]\right)-\Phi(z)\\ &=P\left(\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-g(\theta)\right)\geq\frac{\sqrt{T}\theta}{\alpha}\left[\left(\frac{z}{\sqrt{T}\theta}+1\right)^{-\alpha}-1\right]\right)-\Phi(z),\end{split}$
where in the last line, we use (11) and $g^{\prime}(\theta)=-\alpha\theta^{-\alpha-1}$. Next, we take the short-hand notion $\bar{\Phi}(z)=1-\Phi(z)$ and
$\nu=\frac{\sqrt{T}\theta}{\alpha}\left[\left(\frac{z}{\sqrt{T}\theta}+1\right)^{-\alpha}-1\right].$
It is clear that
$\left\{\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-g(\theta)\right)\geq\nu\right\}=\left\{\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-\mathbb{E}[X]\right)\geq\nu-\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(\mathbb{E}[X]-g(\theta)\right)\right\}.$
Denote $\bar{\Phi}(z)=1-\Phi(z)$. Hence, the triangle inequality implies that
$\begin{split}\left|A(z)\right|&\leq\left|P\left(\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-\mathbb{E}[X]\right)\geq\nu-\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(\mathbb{E}[X]-g(\theta)\right)\right)-\bar{\Phi}\left(\nu-\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(\mathbb{E}[X]-g(\theta)\right)\right)\right|\\ &\quad+\left|\bar{\Phi}\left(\nu-\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(\mathbb{E}[X]-g(\theta)\right)\right)-\bar{\Phi}\left(\nu\right)\right|+\left|\bar{\Phi}\left(\nu\right)-\Phi(z)\right|\\ &\leq d_{Kol}(\sqrt{\frac{T}{\sigma_{2}^{2}}}\left(X-\mathbb{E}[X]\right),\,\xi)+\sqrt{\frac{T}{\sigma_{2}^{2}}}\left|\mathbb{E}[X]-g(\theta)\right|+\left|\bar{\Phi}\left(\nu\right)-\Phi(z)\right|\\ &\leq C\Bigg(d_{Kol}(\sqrt{T}(X-\mathbb{E}[X]),\eta)+\sqrt{T}\left|\mathbb{E}[X]-g(\theta)\right|+\frac{1}{\sqrt{T}}\Bigg),\end{split}$ (12)
where in the last two inequalities we use the standard estimate for the tail of a normal random variable, $\left|\bar{\Phi}(z_{1})-\bar{\Phi}(z_{2})\right|\leq\left|z_{1}-z_{2}\right|$, and Lemma 5.4 of [9]. Subscribing the inequality (12) into (10) yields (10).
The following it taken from Lemma 5.4 of [9].
Lemma 5.4.
Let $c>0$ be a constant. Denote $\nu(z)=\frac{c}{2\beta}\sqrt{T}\Big[\big(1+\frac{z}{c\sqrt{T}}\big)^{-2\beta}-1\Big],\,0<\beta<1$, when $z>-c\sqrt{T}$ and $\bar{\Phi}(z)=1-P(Z\leq z)$. Then there exists some positive number $C$ independent of $T$ such that
$\sup_{z>-c\sqrt{T}}\left|\bar{\Phi}(\nu)-\Phi(z)\right|\leq\frac{C}{\sqrt{T}}.$
Proof 5.5.
We follow the line of the proof of Theorem 3.2 in [24]. By the mean value theorem, there exists some number $\eta(z)\in(0,1)$ such that
$\nu=-z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}.$
Hence,
$\left|\bar{\Phi}(\nu)-\Phi(z)\right|$ $=\left|{\Phi}\Big(\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}\cdot z\Big)-\Phi(z)\right|$
$=\frac{1}{\sqrt{2\pi}}{\int^{z}_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}\,.$
When $z\in(-c\sqrt{T},-\frac{1}{2}c\sqrt{T}]$, it is obvious that
$\frac{1}{\sqrt{2\pi}}{\int_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{z}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}\leq\Phi(-\frac{1}{2}c\sqrt{T})\leq\frac{C}{\sqrt{T}}.$
When $z\in(-\frac{1}{2}c\sqrt{T},\,0]$, we have
$\frac{1}{\sqrt{2\pi}}{\int_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{z}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}$ $\leq\left|z\right|e^{-\frac{z^{2}}{2}}\Big(\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}-1\Big).$ (13)
The mean value theorem implies that there exists some number $\eta^{\prime}\in(0,1)$ such that
$\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}-1=(-1-2\beta)\frac{z\eta}{c\sqrt{T}}\big(1+\frac{z\eta\eta^{\prime}}{c\sqrt{T}}\big)^{-2\beta-2}\leq(1+2\beta)\left(\frac{1}{2}\right)^{-2\beta-2}\frac{\left|z\right|}{c\sqrt{T}}.$
Substituting the above inequality into (13), and since the function $f(z)={z}^{2}e^{-\frac{z^{2}}{2}}$ is uniformly bounded, we have that when $z\in(-\frac{1}{2}c\sqrt{T},\,0]$,
$\frac{1}{\sqrt{2\pi}}{\int_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{z}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}$ $\leq\frac{C}{\sqrt{T}}.$
Denote $a=\sup\left\{z^{2}e^{-\frac{z^{2}}{2}}:\,z\in\mathbb{R}\right\}$ and denote the domain
$D:=\left\{(s,z)\in\mathbb{R}_{+}^{2}:\,\frac{1}{z}\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}\leq s\leq\frac{1}{z},\,z\in(0,\,c\sqrt{T})\right\}.$
We claim that the function
$f_{2}(s,z)=z^{2}e^{-\frac{s^{2}z^{4}}{2}},\quad(s,z)\in D$
is uniformly bounded in the domain $D$. In fact, for the above positive constant $a$, we have
$f_{2}(s,z)=\frac{1}{s^{2}z^{2}}\left((sz^{2})^{2}e^{-\frac{(sz^{2})^{2}}{2}}\right)\leq\frac{a}{(sz)^{2}}\leq a\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{2(2\beta+1)}\leq 2^{2(2\beta+1)}a.$
When $z\in(0,\,c\sqrt{T})$, using the mean value theorem and making the change of variable $t=z^{2}s$ together with the fact that $f_{2}(s,z)$ is uniformly bounded in the above domain $D$, we conclude that there exists a number $\eta^{\prime}\in(0,1)$ such that
${\int_{z\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{z}\,e^{-\frac{t^{2}}{2}}\mathrm{d}t}$ $={\int_{\frac{1}{z}\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}}^{\frac{1}{z}}\,\,z^{2}e^{-\frac{s^{2}z^{4}}{2}}\mathrm{d}s}$
$\leq C\frac{1}{z}\Big(1-\big(1+\frac{z\eta}{c\sqrt{T}}\big)^{-2\beta-1}\Big)$
$=C(1+2\beta)\frac{1}{z}\big(1+\frac{z\eta\eta^{\prime}}{c\sqrt{T}}\big)^{-2\beta-2}\frac{z\eta}{c\sqrt{T}}$
$\leq\frac{C}{\sqrt{T}}.$
Finally, when $z\in[c\sqrt{T},\,\infty)$, we have that
${\Phi}(-z)\leq{\Phi}(-c\sqrt{T})\leq\frac{1}{2c\sqrt{T}}$
and $\nu(z)\leq\frac{c}{2\beta}\sqrt{T}\Big[2^{-2\beta}-1\Big]$. Hence,
$\Phi(\nu)$ $\leq\Phi(\frac{c}{2\beta}\sqrt{T}[2^{-2\beta}-1])\leq\frac{\beta}{[1-2^{-2\beta}]c\sqrt{T}},$
$\left|\bar{\Phi}(\nu)-\Phi(z)\right|$ $=\left|\bar{\Phi}(z)-\Phi(v)\right|\leq\bar{\Phi}(z)+\Phi(\nu)={\Phi}(-z)+\Phi(\nu)\leq\frac{C}{\sqrt{T}}.$
Acknowledgments
The authors wish to express their sincere thanks to the anonymous referees for their valuable suggestions and comments. Y. Chen is supported by the National Natural Science Foundation of China (No. 12461029); Z. Tang are supported by the Project of Basic Research of Science and Technology Plan of Gansu Province in 2025–Excellent doctoral program (25JRRA235) and the “Innovation Star” project of university graduate students in Gansu Province (2025CXZX-867); Y. Li are supported by the Youth Project of Hunan Provincial Natural Science Foundation of China (No. 2024JJ6417), the Project of Education Department of Hunan Province (24B0187) and the 111 Project (No. D23017); H. Yi is supported by the High level Talent Research Project of Liupanshui Normal University (LPSSYKYJJ202416).